Exercise 3.27 Novak produces one single product, a small reading tablet, and sells it at $90 per unit. Its current annual sales are $162,000. Its annual fixed costs include factory rent, $30,780; depreciation expense; equipment, $8,100; utilities, $16,200; insurance, $6,480. Its variable costs include materials, $27 per unit, and direct labour, $36 per unit. Novaks income tax rate is 20%. What is the contribution margin per unit? Contribution margin per unit s LINK TO TEXT LINK TO TEXT LINK TO TEXT What is the contribution margin ratio? Contribution margin ratio LINK TO TEXT LINK TO TEXT LINK TO TEXT How many units must Novak sell to break even? (Round answer to the next whole integer, e.g. 1,255.) Sales units LINK TO TEXT LINK TO TEXT LINK TO TEXTIf Novak would like to earn a profit after tax of $14,000, what should the sales be? At this sales level, what is the degree of operating leverage? What is the margin of safety in unit? (Round total sales and margin of safety to O decimal places, e.g. 5,125. Round degree of operating leverage to 2 decimal places, e.g. 15.16.) Total Sales Degree of operating leverage Margin of safety units LINK TO TEXT LINK TO TEXTLINK TO TEXT If Novak would like to earn a profit after tax that is 8% of sales, what should the sales be? How many units does Novak need to increase from the current sales level? Total Sales Units to be increased by units LINK TO TEXT LINK TO TEXT LINK TO TEXT
Answer #1
Req 1.
Contribution margin per unit:
Selling price per unit 90
Less: Variable cost per unit
Material cost per unit 27
labour cost per unit 36
Total Variable cost per unit 63
Contribution margin per unit: 27
Req 2.
Contribution margin ratio:
Contribution margin per unit 27
Divide: Selling price 90
CM ratio 30%
Req 3.
Total fixed cost:
Factory rent 30780
Ddepreciation 8100
Utilities 16200
Insurance 6480
Total fixed cost: 61560
Divide: Contribution margin per unit 27
Break even units 2280
Req 4.
After tax profits 14000
tax rate 20%
Desired profits after tax 17500
(14000/80 *100)
Add: Fixed cost 61560
Desired contribution 79060
Divide: CM ratio 30%
Desired sales in $ 263533
Desired sales in Units 2928 units
(263533 / 90)
Degree of Operating leverage:
Contribution margin (2928*27) 79060
Divide: Net Income before tax 17500
Degree of Operating leverage: 4.52
Margin of Safety:
Sales required 2928 units
Less: Break even units 2280 units
Margin of Safety: 648 units
Req 5.
After tax Profits as % of sales 8%
Tax rate 20%
Before tax desired profits as % of sales 10%
(8% /80% *100)
Variable cost ratio as % of sales 70%
(100- CM ratio i.e. 30%)
Combined of profits and variable cost 80% of sales
Fixed cost as % of sales should be 20% of sales
Now, Total fixed cost = 61560
It should be 20% of sales
Therefore, Desired sales = 61560 /20% = $ 307800
Desired sales in units = 307800/90 = 3420 units